In the amusement park ride Mr. Freeze at Six Flags, riders are uniformly accelerated from rest by magnetic induction motors along a 70 meter horizontal track in just 5 seconds. While accelerating, friction and air drag exert 500N of force on the train. The train then coasts through the loops and turns for the remainder of the ride. A typical train loaded with passengers has a mass of 2500 kg.



1
Expert's answer
2021-05-14T09:41:59-0400

Displacement, "s=70\space m"

Time, "t=5\space s"

Initial velocity, "u=0"

"s=ut+\dfrac{1}{2}at^2"

"70=\dfrac{1}{2}(a\times25)"

"a=\dfrac{140}{25}=5.6\space m/s^2"

Resisting forces "=500\space N"

Resultant force in forward direction, "F'=ma-Resisting \space \space forces"

"F'=(2500\times5.6)-500"

"F'=13500\space N"